What is the equation of the line that goes through the points (-1, 3) and (4, -7) in general form?
Answer by Cromlix(4381)


Gradient = y2 - y1/ x2 - x1 => -7 - 3/4 - (-1)
= - 10/5 = -2
Using y - b = m(x - a) and coords (4, -7)
y - (-7) = -2(x - 4)
y + 7 = -2x + 8
y = -2x + 8 - 7
y = -2x + 1